OpenStudy (anonymous):

Find the distance (to the nearest km) from: X (20ºS, 86ºW) to Y (50ºN, 86ºW). If you could please help me with this question, it would be greatly appreciated.

OpenStudy (anonymous):

Lat= 50 +20 = 70 N.Mile= 70 * 60 =130 KM= 130 * 1.852 = 240.76 KM